OpenStudy (anonymous):

What is the mean of the following data without the outlier? 41, 44, 50, 105 52 45 60 33.75 @lupita1995

OpenStudy (anonymous):

60

OpenStudy (anonymous):

I think it's 60.

OpenStudy (solomonzelman):

I almost thought that there should be a pattern. Would be hard to find one :)

OpenStudy (anonymous):

The mean is 41+44+50+52+60+33.75+45 all divided by 7 (remember to discard the outlier). 85+102+105+33.75)/7=202+90+33.75)/7=295.75/7..that's 40+2+1.75/7..42+1/7+3/28......42+7/28=42 1/4

OpenStudy (anonymous):

So the answer would be?

OpenStudy (anonymous):

Is it 60?

OpenStudy (anonymous):

it would say so

OpenStudy (anonymous):

60